# Drop your Decoder: Pre-training with Bag-of-Word Prediction for Dense Passage Retrieval

### v1

@@ 3.3. Pre-training with Bag-of-Word Prediction
Building upon the above observations about token compositions and domain adaption abilities, here we propose a new pre-training schema that explicitly removes the mandatory needed decoder in existing MAE-style pre-training. As is shown in Algorithm 2, the sentence representations at the [CLS] token position are first projected to the vocabulary space with the BERT language-model heads (i.e., its embedding matrix $E^{T}$). Then we apply a multi-label cross-entropy loss to predict the input token set $T$ directly. Bag-of-Word means the label $T$ is a set of tokenized input text. For simplicity, it is not related to positional information, similar to the traits of sparse retrieval (Robertson et al., 2009).
@@
There are several benefits of the design behind our method. First of all, Bag-of-Word prediction is a direct compression of input lexicon information into dense representations with extreme simplicity, which merely utilizes the existing BERT language model head for projection. Secondly, it removes the need for extra Transformers-based decoder layers, data preprocessing, or extensive data masking, which introduces ZERO extra computational complexity, compared to other expensive MAE-style pre-training. Moreover, different from previous MAE-style pre-training modifications, which mostly focus on extensive modifications to the decoder tasks (Liu and Shao, 2022; Wang et al., 2022; Wu et al., 2023; Zhou et al., 2022), Bag-of-Word pre-training only involves a simple lexicon prediction task, which bridges the condensed representations and lexicon traits. Such a pre-training schema is highly interpretable and easy to implement.
Additional MS-MARCO Natural Question Trivia QA
Model Decoder MRR@10 R@50 R@1k R@5 R@20 R@100 R@5 R@20 R@100
BM25 (Robertson et al., 2009) 18.7 59.2 85.7 - 59.1 73.7 - 66.9 76.7
SPLADE (Formal et al., 2021) 32.2 - 95.5 - - - - - -
DPR (Karpukhin et al., 2020) 31.7 58.0 85.7 - 74.4 85.3 - 79.3 84.9
ANCE (Xiong et al., 2020) 33.0 - 95.9 - 81.9 87.5 - 80.3 85.3
SEED (Lu et al., 2021) $\checkmark$ 33.9 - 96.1 - 83.1 88.7 - - -
RocketQA (Qu et al., 2021) 37.0 - 97.9 74.0 82.7 88.5 - - -
SimLM (Wang et al., 2022) $\checkmark$ 39.1 87.3 98.6 73.6 84.3 89.0 78.1 83.8 87.6
RetroMAE (Liu and Shao, 2022) $\checkmark$ 39.3 87.0 98.5 74.4 84.4 89.4 78.9 84.5 88.0
BoW Prediction 40.1* 88.7* 98.9 75.3 84.6 90.4* 79.4* 84.9* 88.0
Data Process Additional Decoder Training Speed
Further Pre-train Encoder $O(n)$ 0.0476 - - 269.708 -
Auto-Encoding (Gao and Callan, 2021, 2022; Liu and Shao, 2022; Wu et al., 2023) Encoder-Decoder $O(n)$ 0.0940 $O(n^{2})$ 0.0013 222.658 17.4%
Enhanced Decoding (Liu and Shao, 2022) Encoder-Decoder $O(n^{2})$ 5.6261 $O(n^{2})$ 0.0012 85.797 68.2%
BoW Prediction Encoder $O(n)$ 0.0533 - 0.0002 266.359 1.2%
Table 4. Comparisons of architecture, data process, additional decoder, and training speed of different pre-training methods. Note that Bag-of-Word (BoW) prediction does not involve additional decoders, thus we report the GPU computation time of the prediction task. The original implementation of enhanced decoding (Liu and Shao, 2022) randomly generates 256 attention masks for each input text, and picks one as the final mask used in the pre-training. And thus it has $O(n^{2})$ data process complexity.
4. Experiments
4.1. Experiment Settings
Pre-training
Wikipedia (Foundation, [n. d.]) and Bookcorpus (Zhu et al., 2015), the original BERT training sets, are used as pre-training corpora, which are commonly adopted in previous retrieval pre-trainings (Gao and Callan, 2021; Liu and Shao, 2022; Lu et al., 2021). The BERT-base-uncased checkpoint is used to initialize the encoder for saving computation. The model is pre-trained using the AdamW optimizer for 20 epochs, with a learning rate of 3e-4, a batch size of 2048, and a maximum sequence length of 512. Following previous in-domain pre-training settings (Gao and Callan, 2021; Liu and Shao, 2022; Lu et al., 2021; Wang et al., 2022), MS-MARCO corpus (Nguyen et al., 2016) is also used when evaluating the MS-MARCO benchmark. All pre-training hyper-parameters are kept the same, except the maximum sequence length is cut to 144.
Evaluation
We conduct supervised evaluations on several large-scale web search and open-domain QA datasets, including MS-MARCO passage ranking (Nguyen et al., 2016), Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017). The metrics of Mean Reciprocal Rank (MRR@k) and Recall (R@k) are reported, following (Gao and Callan, 2022). We fine-tune MS-MARCO with a batch size of 64, a learning rate of 2e-5, and a passage group size of 16 for 3 epochs. Following (Karpukhin et al., 2020), we utilize the DPR-Wiki versions of NQ and TriviaQA benchmarks and evaluate retrieval performances. We fine-tuned NQ and TriviaQA with the same hyper-parameters, except that the learning rate is 5e-6, and the number of training epochs is 20. Retrievers are fine-tuned with ANCE manners (Xiong et al., 2020) by following (Gao and Callan, 2021, 2022; Liu and Shao, 2022). Following (Gao and Callan, 2021; Liu and Shao, 2022), both BM25 negatives and retriever-mined negatives are used in fine-tuning stages for fair comparisons. A large-scale heterogeneous zero-shot benchmark BEIR (Thakur et al., 2021) is also used to evaluate the out-of-domain transferability of our method. Normalized Discounted Cumulative Gain (NDCG@10) is used as the metric for BEIR.
@@ 4.1. Experiment Settings
BoW Prediction 40.1 88.7 98.9 75.3 84.6 90.4 79.4 84.9 88.0
Table 5. Retrieval performances of different pre-training schema with BM25 negatives and mined negatives.
@@
We include multiple public sparse and dense retrieval baselines for comparisons. Sparse retrieval baselines include BM25 (Robertson et al., 2009) and SPLADE (Formal et al., 2021). Dense retrieval baselines are composed of multiple well-known pre-training designs, including MAE-style pre-training SEED (Lu et al., 2021), Condenser (Gao and Callan, 2021), coCondenser (Gao and Callan, 2022), SimLM (Wang et al., 2022), RetroMAE (Liu and Shao, 2022), and other dense retrieval modifications, like DPR (Karpukhin et al., 2020), ANCE (Xiong et al., 2020), RocketQA (Qu et al., 2021). Particularly, we reproduce multiple MAE-style pre-training methods and compare with their performances and training efficiencies, including Auto-Encoding (Liu and Shao, 2022), Auto-Regression (Lu et al., 2021), and Enhanced Decoding (Liu and Shao, 2022).
Other data augmentation techniques, such as context-based (Wu et al., 2023), expansion-based (Wu et al., 2022), mutual information-based (Li et al., 2023) or multi-task (Zhou et al., 2022) pre-training, are not included here for fair comparisons. Our proposed Bag-of-Word prediction is orthogonal to them, and we will leave the combinations of our method with various augmentation techniques to further work.
@@ 4.2. Main Results
Bag-of-Word prediction pre-trains dense retrieval models with higher speed and fewer CPU/GPU complexities. As is shown in Table 4, compared with MAE-based encoder-decoder structures, our method removes the need for the additional decoder, which saves ¿ 83% GPU computation time on the decoding side. Additionally, our method does not involve extensive masking strategies, which saves 99% CPU time for the data process compared to enhanced decoding. Overall, the Bag-of-Word prediction has a very slight training speed degeneration and achieves 67%, 19%, and 16.2% training speedup compared with enhanced decoding, auto-regression, and auto-encoding.
@@
Table 6. Ablation study on the heterogeneous BEIR benchmark for our Bag-of-Word prediction pre-training method, with various encoder mask ratios or removing the proposed Bag-of-Word prediction task.
@@ 5.1. Performances with Different Negatives
We compare the retrieval performances of different pre-training schemas and report the results with different negative training data in Table 5. The retrievers are trained with BM25 negatives and retriever-mined negatives, respectively, by following previous supervised fine-tuning pipelines (Xiong et al., 2020; Gao and Callan, 2021). It can be seen that even with different negatives, Bag-of-Word prediction pre-training surpasses all MAE-style pre-training, MLM further pre-training, and BERT baselines, under the same pre-training and fine-tuning pipelines. That further verifies our pre-training method is simple but more effective than other MAE pre-training schemas.
@@ 5.3. Transferability and Encoder Mask Ratios
Previous study (Liu and Shao, 2022; Wu et al., 2023) reported that properly increasing the mask ratio for the encoder can lead to higher retrieval performances. Hence we perform pre-training with Bag-of-Word prediction under different encoder mask ratios and report the performances on the heterogeneous BEIR benchmarks. BEIR benchmarks include 18 out-of-domain retrieval datasets across different domains, sizes, and topics, which is suitable for evaluating the transferability of retriever models. As is shown in Table 6, pre-training with Bag-of-Word prediction achieves relatively +6.39% performance gains over the BERT baseline. On the contrary, removing the Bag-of-Word prediction task and applying further pre-training with pure MLM loss will incur +1.89% (6.39% $\rightarrow$ 4.50%) performance degeneration. Here we also observe that increasing the encoder mask ratio from 15% to 30% or 45% does not yield higher retrieval performances.
@@ 6. Conclusion
In this study, we explore the direct application of Bag-of-Word prediction in dense retrieval pre-training. There are also effective modifications about data augmentation or task-ensembling techniques, such as context sampling (Wu et al., 2023), document expansion (Nogueira et al., 2019; Wu et al., 2023), mutual information masking (Li et al., 2023) and multi-task (Zhou et al., 2022). Bag-of-Word prediction focuses on pre-training paradigm innovation, which is orthogonal to them. Combinations of Bag-of-Word prediction with those data augmentation or task-ensembling techniques will be likely to further boost the downstream retrieval performances, and we leave this to our future work.

### v2

@@ 3.3. Pre-training with Bag-of-Word Prediction
Building upon the above observations about token compositions and domain adaptation abilities, here we propose a new pre-training schema that explicitly removes the mandatory needed decoder in existing MAE-style pre-training. As is shown in Algorithm 2, the sentence representations at the [CLS] token position are first projected to the vocabulary space with the BERT language-model heads (i.e., its embedding matrix $E^{T}$). Then we apply a multi-label cross-entropy loss to predict the input token set $T$ directly. Bag-of-Word means the label $T$ is a set of tokenized input text. For simplicity, it is not related to positional information, similar to the traits of sparse retrieval (Robertson et al., 2009).
@@
BoW Prediction 40.1* 88.7* 98.9 75.3* 84.6 90.4 79.4* 84.9 88.0
Table 5. Retrieval performances of different pre-training schema with BM25 negatives and mined negatives. * means significant improvement over the best results of reproduced MAE pre-training baselines. ($p\leq 0.05$)
There are several benefits of the design behind our method. First of all, Bag-of-Word prediction is a direct compression of input lexicon information into dense representations with extreme simplicity, which merely utilizes the existing BERT language model head for projection. Secondly, it removes the need for extra Transformers-based decoder layers, data preprocessing, or extensive data masking, which introduces ZERO extra computational complexity, compared to other expensive MAE-style pre-training. Moreover, different from previous MAE-style pre-training modifications, which mostly focus on extensive modifications to the decoder tasks (Xiao et al., 2022; Wang et al., 2023b; Wu et al., 2023a; Zhou et al., 2022), Bag-of-Word pre-training only involves a simple lexicon prediction task, which bridges the condensed representations and lexicon traits. Such a pre-training schema is highly interpretable and easy to implement.
4. Experiments
4.1. Experiment Settings
Pre-training
Wikipedia (Foundation, [n. d.]) and Bookcorpus (Zhu et al., 2015), the original BERT training sets, are used as pre-training corpora, which are commonly adopted in previous retrieval pre-trainings (Gao and Callan, 2021; Xiao et al., 2022; Lu et al., 2021). The BERT-base-uncased checkpoint is used to initialize the encoder for saving computation. The model is pre-trained using the AdamW optimizer for 20 epochs, with a learning rate of 3e-4, a batch size of 2048, and a maximum sequence length of 512. Following previous in-domain pre-training settings (Gao and Callan, 2021; Xiao et al., 2022; Lu et al., 2021; Wang et al., 2023b; Yu et al., 2024), MS-MARCO corpus (Nguyen et al., 2016) is also used when evaluating the MS-MARCO benchmark for making fair comparisons with existing works (Gao and Callan, 2021; Xiao et al., 2022; Wang et al., 2023b). All pre-training hyper-parameters are kept the same, except the maximum sequence length is cut to 144.
Evaluation
We conduct supervised evaluations on several large-scale web search and open-domain QA datasets, including MS-MARCO passage ranking (Nguyen et al., 2016), Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017). The metrics of Mean Reciprocal Rank (MRR@k) and Recall (R@k) are reported by following (Gao and Callan, 2022). We fine-tune MS-MARCO with a batch size of 64, a learning rate of 2e-5, and a negative size of 15 for 3 epochs. Following (Karpukhin et al., 2020), we utilize the DPR-Wiki versions of NQ and TriviaQA benchmarks and evaluate retrieval performances. We fine-tuned NQ and TriviaQA with the same hyper-parameters, except that the learning rate is 5e-6, and the number of training epochs is 20. Retrievers are fine-tuned with ANCE manners (Xiong et al., 2021) by following (Gao and Callan, 2021, 2022; Xiao et al., 2022). Following (Gao and Callan, 2021; Xiao et al., 2022), both BM25 negatives and retriever-mined negatives from previous stage are used in fine-tuning stages for fair comparisons. A large-scale heterogeneous zero-shot benchmark BEIR (Thakur et al., 2021) is also used to evaluate the out-of-domain transferability of our method. Normalized Discounted Cumulative Gain (NDCG@10) is used as the metric for BEIR.
@@ 4.1. Experiment Settings
We include multiple public sparse and dense retrieval baselines for comparisons. Sparse retrieval baselines include BM25 (Robertson et al., 2009) and SPLADE++ (Formal et al., 2022). Dense retrieval baselines are composed of multiple well-known pre-training designs, including MAE-style pre-training SEED (Lu et al., 2021), Condenser (Gao and Callan, 2021), coCondenser (Gao and Callan, 2022), SimLM (Wang et al., 2023b), RetroMAE (Xiao et al., 2022), and other dense retrieval modifications, like DPR (Karpukhin et al., 2020), ANCE (Xiong et al., 2021), RocketQA (Qu et al., 2021). Distillations from re-rankers are not included as their retrieval performances rely more on the capability of re-ranker teachers, while in our study we focus on inspecting the dense retrieval per-training of retrievers. Particularly, we reproduce multiple MAE-style pre-training methods with same hyperparameters and compare with their performances and training efficiencies, including Auto-Encoding (Xiao et al., 2022), Auto-Regression (Lu et al., 2021), and Enhanced Decoding (Xiao et al., 2022).
Other data augmentation techniques, such as context-based (Wu et al., 2023a), expansion-based (Wu et al., 2023b), mutual information-based (Li et al., 2023) or multi-task (Zhou et al., 2022) pre-training, are not included here for fair comparisons. Our proposed Bag-of-Word prediction is orthogonal to them, and we will leave the combinations of our method with various augmentation techniques to further work. Hybrid retrievals (Liu et al., 2023; Wu et al., 2023c) are not included in the baselines, as our work only focuses on the pre-training for dense retrieval.
@@ 4.2. Main Results
Bag-of-Word prediction pre-trains dense retrieval models with higher speed and fewer CPU/GPU complexities. We compare the training efficiencies with a single NVIDIA H800 GPU. As is shown in Table 4, compared with MAE-based encoder-decoder structures, our method removes the need for the additional decoder, which saves ¿ 83% GPU computation time on the decoding side. Additionally, our method does not involve extensive masking strategies, which saves 99% CPU time for the data process compared to enhanced decoding. Overall, the Bag-of-Word prediction has a very slight training speed degeneration and achieves 67%, 19%, and 16.2% training speedup compared with enhanced decoding, auto-regression, and auto-encoding.
@@
Table 6. Ablation study on the heterogeneous BEIR benchmark for our Bag-of-Word prediction pre-training method, with various encoder mask ratios or removing the proposed Bag-of-Word prediction task. * means significant improvements over the wo/BoW settting. ($p\leq 0.05$)
@@ 5.1. Performances with Different Negatives
We compare the retrieval performances of different pre-training schemas and report the results with different negative training data in Table 5. The retrievers are trained with BM25 negatives and retriever-mined negatives, respectively, by following previous supervised fine-tuning pipelines (Xiong et al., 2021; Gao and Callan, 2021). It can be seen that even with different negatives, Bag-of-Word prediction pre-training surpasses all MAE-style pre-training, MLM further pre-training, and BERT baselines, under the same pre-training and fine-tuning pipelines. This further verifies that our pre-training method is simple but more effective than other MAE pre-training schemas.
@@ 5.3. Transferability and Encoder Mask Ratios
Previous study (Xiao et al., 2022; Wu et al., 2023a) reported that properly increasing the mask ratio for the encoder can lead to higher retrieval performances. Hence we perform pre-training with Bag-of-Word prediction under different encoder mask ratios and report the performances on the heterogeneous BEIR benchmarks. BEIR benchmarks include 18 out-of-domain retrieval datasets across different domains, sizes, and topics, which is suitable for evaluating the transferability of retriever models. As is shown in Table 6, pre-training with Bag-of-Word prediction achieves relatively +6.39% performance gains over the BERT baseline. On the contrary, removing the Bag-of-Word prediction task and applying further pre-training with pure MLM loss will incur 1.89% (6.39% $\rightarrow$ 4.50%) performance degeneration. Here we also observe that increasing the encoder mask ratio from 15% to 30% or 45% does not yield higher retrieval performances.
@@ 6. Conclusion
In this study, we explore the direct application of Bag-of-Word prediction in dense retrieval pre-training. There are also effective modifications about data augmentation or task-ensembling techniques, such as context sampling (Wu et al., 2023a), document expansion (Nogueira et al., 2019; Wu et al., 2023a), mutual information masking (Li et al., 2023) and multi-task (Zhou et al., 2022). Bag-of-Word prediction focuses on pre-training paradigm innovation, which is orthogonal to them. Combinations of Bag-of-Word prediction with those data augmentation or task-ensembling techniques will be likely to further boost the downstream retrieval performances, and we leave this to our future work.